EVIDENCE OF RESONANT COLLISIONS IN VELOCITY SELECTIVE STUDIES OF THE DIPOLE-DIPOLE COLLISION CROSS SECTION OF ^{13}CH_{3}F

\underline{LIAM M. DUFFY}, Department of Physics, Duke University, Durham, NC 27706; DAVID SKA-TRUD, Army Research Office, P.O. Box 12211, RTP, NC 27709.

For molecules with permanent electric dipole moments, the most rapid energy transfer mechanism is due to dipoledipole collisions which induce ΔJ = +/- 1 transitions. Although this process is the dominant contributor to pressure broadening, little is known about the temperature dependence of the process. We have used a velocity selective time resolved pump/probe technique based on infrared-millimeter/sub-millimeter wave double resonance spectroscopy to study the temperature dependence of the dipole-dipole cross section of ^{13}CH_{3}F. In these experiments we extract the collision cross section as a function of mean relative velocity without changing the rotational distribution of the collision partners. Our results show that the dipole-dipole collision cross section is largely independent of the mean relative velocity in the range between 400 and 900 m/s. On the other hand, we clearly observe a decrease in the cross section when the gas cell temperature is raised from 173K to 400K. These results strongly suggest that resonant collisions dominate the dipoledipole collision cross section. We will discuss these finds in terms of the temperature scaling predicted by Anderson, Tsao and Curnutte theory.